OpenStudy (anonymous):

Solve for m: 2/3 = m/42

OpenStudy (anonymous):

m = 42*2/3 m = 84/3 m = 28

Nnesha (nnesha):

multiply 42 both side

Nnesha (nnesha):

 because its dividing on right side to move we have to multiply